OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 20x2 – 28x – 48 and check your factors for accuracy.

OpenStudy (anonymous):

first divide by the gcf of the numbers , 4 = 4(5x^2 - 7x - 12) = 4(5x -12)(x + 1)

OpenStudy (anonymous):

factorising the trinomial - 5x and x must be first term in each bracket. Then you need to find two numbers whcih when multiplied together must equal -12 and when combined with 5 and then added equal - 7.And -12 and +1 are those numbers.

OpenStudy (anonymous):

hope that helps. they are a bit tricky especially when the number before the x^2 is greater than 1.